# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 3.4.3 Process Supervision
In the post-training stage, the focus of supervision shifts from the correctness of the outcome to the rationality of the process. Datasets with Chain-of-Thought annotations, such as Visual CoT(shao2024visual) and LLaVA-CoT-100K(xu2024llava), provide the model with detailed reasoning paths, enabling it to perform efficient behavioral cloning through supervised fine-tuning or to undergo goal-oriented optimization based on reward signals via reinforcement learning methods.
During the inference stage, the supervision mechanism manifests in a more dynamic and immediate manner. Chain-of-Thought prompting(wei2022chain; kojima2022large) acts as a non-parametric, on-the-fly form of soft supervision, efficiently guiding the model into a structured reasoning mode. More advanced guidance mechanisms, such as the Process Reward Model proposed by VisVM(wang2024scaling), or the ”critic” role in the multi-agent framework employed by Elhenawy et al.(elhenawy2024visual), constitute external dynamic supervisory signals at inference time. Meanwhile, self-optimization mechanisms, as represented by R3V(cheng2024vision) and EVLM(khalid2024evlm), embody a higher-order form of internal dynamic supervision. In this mechanism, the model acts as its own critic, forming an effective closed feedback loop by reflecting on, reviewing, and generating corrective instructions for the intermediate steps it has already produced. Whether originating externally or internally, these dynamic supervisory signals collectively provide fine-grained, real-time calibration of the model’s reasoning behavior.
@@ 4.1 Mathematical and Logical Reasoning
Benchmarks such as MathVista(lu2023mathvista), MATH-Vision(wang2024measuring), We-Math(qiao2024we), and CMM-Math(liu2024cmm) integrate visual information into mathematical reasoning tasks. MathVerse(zhang2024mathverse) aims to evaluate models’ reasoning abilities on math problems involving charts and adopts a chain-of-thought evaluation strategy, leveraging GPT-4(V) for fine-grained assessment of generated answers. ErrorRadar(yan2024errorradar) introduces a new multimodal error detection task in a formalized manner, which includes two subtasks: error step identification and error categorization, thus providing a comprehensive framework for assessing complex mathematical reasoning abilities in multimodal large language models.
Additionally, LogicVista(xiao2024logicvista) targets the evaluation of comprehensive logical reasoning in visual environments, while MME-Reasoning(yuan2025mme) offers a thorough assessment of logical reasoning capabilities, encompassing inductive, deductive, and abductive reasoning. EnigmaEval(wang2025enigmaeval) explores models’ abilities to integrate implicit knowledge and perform multi-step deductive reasoning by presenting puzzles of varying complexity. MuCR(li2024multimodal), a novel multimodal causal reasoning benchmark, leverages synthetic twin image-text pairs to challenge models in identifying causal relationships across different modalities.
@@ 4.2 Spatiotemporal and Directional Reasoning
PulseCheck457(wang2025pulsecheck457) offers a comprehensive framework for assessing 6D spatial reasoning across varying levels of complexity. It includes seven question types and spans five difficulty levels, ranging from basic single-object recognition to advanced 6D spatial reasoning tasks. GSR-BENCH(rajabi2024gsr) extends the What’sUp dataset and proposes a novel integrated evaluation method for understanding spatial relations. CDR(yin2025multimodal) targets the assessment of directional reasoning in multimodal large language models. It includes images annotated with spatial directions (e.g., up, down, left, right) and compass directions (e.g., north, south, east, west), thereby addressing a previously underexplored area in directional reasoning.For temporal reasoning, TOMATO(shangguan2024tomato) introduces a new benchmark to assess models’ ability to reason about temporal dynamics in video understanding. In the context of autonomous driving, DriveLMM-o1(ishaq2025drivelmm) provides a benchmark with over 4,000 visual question-answering instances, covering perception, prediction, and planning tasks. Each instance includes a detailed step-by-step reasoning process.
@@ 4.3 Vision-Language Transformation Reasoning
HumanEval-V(zhang2024humaneval) employs a novel code generation task to comprehensively evaluate a model’s chart comprehension capabilities. It comprises six task types, each featuring carefully curated charts along with function signatures and test cases. Plot2Code(wu2024plot2code) provides a comprehensive visual-to-code benchmark that includes high-quality matplotlib charts collected from the matplotlib gallery, accompanied by the corresponding source code and descriptive instructions summarized by GPT-4. Similarly, Plot2XML(cui2025draw) offers a benchmark for reconstructing scientific plots into editable XML code. P2GB(chen2024plug) quantitatively evaluates the visual reasoning abilities of multimodal large language models in rich-text or high-resolution image contexts. Current benchmarks for multimodal chain-of-thought (CoT) reasoning predominantly follow the conventional paradigm of multimodal input paired with textual output. To address this limitation, Cheng et al.(cheng2025comt) proposed CoMT, which requires both multimodal input and multimodal reasoning output, aiming to emulate human behavior in integrating visual operations during the reasoning process.
@@ 4.4 Sequential and Multi-Image Reasoning
MIMU(zhang2025benchmark) is developed to assess the performance of multimodal large language models in multi-image input scenarios. It focuses on the model’s ability to perceive fine-grained visual details, with an emphasis on two specific tasks: image-to-image matching and multi-image-to-text matching.
@@ 4.5 Multimodal Integrated Reasoning
MME-CoT(jiang2025mme), the first comprehensive benchmark for evaluating the CoT reasoning capabilities of multimodal large language models, spans six domains: mathematics, science, optical character recognition (OCR), logic, spatiotemporal reasoning, and general scenarios. M3CoT(chen2024m3cot) takes the first step toward addressing multi-domain, multi-step, and multimodal scenarios in multimodal CoT reasoning. Similarly, EMMA(hao2025can) and VRC-Bench[84] offer multi-step benchmarks that assess CoT reasoning across modalities and domains. In addition, R1-Onevision-Bench(yang2025r1) spans multiple subjects—including mathematics, physics, chemistry, biology, and logical reasoning—covering examination content from middle school to university and beyond. NPHardEval4V(fan2024nphardeval4v) serves as a dynamic benchmark designed to isolate the effects of various factors such as image recognition and instruction following from overall model performance, thereby focusing the evaluation on reasoning capability. Likewise, VisualPuzzles(song2025visualpuzzles) reduces the reliance on domain-specific knowledge while increasing reasoning complexity to better assess a model’s multimodal reasoning performance.
@@ 4.6 Reasoning Process Quality Evaluation
MPBench(xu2025mpbench) is a comprehensive multi-task benchmark that systematically evaluates the effectiveness of Process Reward Models (PRMs) through three evaluation paradigms: step correctness identification, answer aggregation, and reasoning path search. MiCEval(zhou2024miceval) assesses the correctness of reasoning chains by evaluating both the quality of model-generated descriptions and each reasoning step. The description evaluation focuses on the accuracy of image descriptions, while the reasoning step evaluation emphasizes the quality of each specific inference. VisualProcessBench(wang2025visualprm), a benchmark with manually annotated step-by-step correctness labels, is primarily used to assess the ability of PRMs to detect incorrect reasoning steps in multimodal reasoning tasks. In addition, MMIR(yan2025multimodal) and the work by Jia et al.(jia2025exploring) focus on identifying inconsistency issues in multimodal reasoning, offering valuable insights for future improvements of multimodal large language models.
@@
MathVista(lu2023mathvista) 2024 Math QA T, I Accuracy MC, Open 6,141
MATH-Vision(wang2024measuring) 2024 Math QA T, I Accuracy MC, Open 3,040
We-Math(qiao2024we) 2024 Math QA T, I Custom Metric MC 6.5K
LogicVista(xiao2024logicvista) 2024 Logical Reasoning T, I Accuracy MC, Open 448
MathVerse(zhang2024mathverse) 2024 Math QA T, I Accuracy, CoT-Evaluation MC, Open 15,672
ErrorRadar(yan2024errorradar) 2024 Error Detection T, I Accuracy Open 2,500
CMM-Math(liu2024cmm) 2024 Math QA T, I Accuracy MC, Open 28,069
MuCR(li2025multimodal) 2025 Causal Reasoning T, I Custom Metric Open ——
EnigmaEval(wang2025enigmaeval) 2025 Puzzle Solving T, I Accuracy Open 1,184
MME-Reasoning(yuan2025mme) 2025 Logical Reasoning T, I Custom Metric MC, Open, Rule-based 1,188
@@
CDR(yin2025multimodal) 2024 Directional Reasoning T, I Accuracy MC 100K
GSR-BENCH(rajabi2024gsr) 2024 Spatial Reasoning T, I Accuracy MC 4,958
TOMATO(shangguan2024tomato) 2024 Temporal Reasoning T, V Custom Metric MC 1,484
PulseCheck457(wang2025pulsecheck457) 2025 Spatial Reasoning T, I RPDR Open ——
DriveLMM-o1(ishaq2025drivelmm) 2025 Autonomous Driving T, I, PC Custom Metric MC, Open ——
@@
Plot2Code(wu2024plot2code) 2024 Visual Coding T, I Custom Metric Open 132
P2G(chen2024plug) 2024 Visual Reasoning T, I Accuracy MC 2,130
HumanEval-V(zhang2024humaneval) 2025 Visual Coding T, I pass@k Open 100K
CoMT(cheng2025comt) 2025 Multi-task T, I Accuracy, Macro-F1 MC 3,853
Plot2XML(cui2025draw) 2025 Visual Coding T, I Custom Standard Open 247
@@
MIMU(zhang2025benchmark) 2025 Multi-task T, I Accuracy Open ——
@@
NPHardEval4V(fan2024nphardeval4v) 2024 VQA T, I Custom Metric Open ——
M3CoT(chen2024m3cot) 2024 VQA T, I Accuracy MC 11,459
EMMA(hao2025can) 2025 VQA T, I pass@k MC, Open 2,788
VRC-Bench(thawakar2025llamav) 2025 Multi-task T, I Custom Multi-metric MC, Open ——
MME-CoT(jiang2025mme) 2025 VQA T, I Quality, Robustness, Efficiency MC, Open 1,130
R1-Onevision-Bench(yang2025r1) 2025 VQA T, I Accuracy MC, Open 942
VisualPuzzles(song2025visualpuzzles) 2025 VQA T, I Accuracy MC 1,168
@@
MiCEval(zhou2024miceval) 2025 VQA T, I Custom Metric Open 2,130
MMIR(yan2025multimodal) 2025 Consistency Reasoning T, I Accuracy MC, Open 534
VisualProcessBench(wang2025visualprm) 2025 Science QA T, I Macro-F1 MC, Open 2,866
4.7 Evaluation Metrics
At present, most existing works(xiang2024atomthink; thawakar2025llamav; li2024multimodal; shangguan2024tomato; wu2024plot2code) assess the reasoning capabilities of LLMs and MLLMs by directly evaluating the accuracy of their final answers, using this as a proxy for chain-of-thought reasoning performance. Alternatively, some studies propose task-specific evaluation metrics. However, these approaches often suffer from limitations such as a lack of comprehensive evaluation and inability to accurately assess complex visual reasoning tasks. To address these issues, MME-CoT(jiang2025mme) proposed the first dedicated evaluation framework for assessing CoT reasoning capabilities in MLLMs, which covers three key aspects: CoT quality, CoT robustness, and CoT efficiency. In addition, MMMR(tie2025mmmr) introduced a modular Reasoning Trace Evaluation Pipeline (RTEP) that provides a holistic assessment of reasoning steps based on three dimensions: relevance to the question, relevance to the answer, and internal consistency across reasoning steps.
@@ 5.1 Embodied AI
In the domain of task decomposition and execution, the works by Hao et al.(hao2025embodied) and Embodied Reasoner(zhang2025embodied) demonstrate the feasibility of automatically dividing operation instructions or task goals into concrete, detailed execution steps. Specifically, Hao et al. enhance humanoid robots’ abilities to perform localized operations in unstructured complex scenarios by incorporating foundation models, while Embodied Reasoner targets interactive search tasks, showcasing the agent’s capacity for autonomous goal recognition, reasoning, and decision-making, thereby strengthening adaptability and proactivity in dynamic tasks.
@@
Regarding task planning and hierarchical coordination, EmbodiedGPT(mu2023embodiedgpt) proposes a unified architecture that integrates high-level strategic planning with low-level control execution, enabling embodied agents to accomplish complex tasks through stepwise planning. The MCoT-Memory framework(liangmemory) introduces a memory retrieval–based, scene graph–driven chain-of-thought mechanism that retains high-confidence experiential knowledge to effectively support long-term and complex task planning in dynamic environments. ECoT(zawalski2024robotic) further expands embodied intelligence’s multimodal reasoning capabilities by incorporating chain-of-thought training mechanisms into Vision-Language-Action (VLA) models, allowing for multi-step task reasoning and subtask decomposition with strong generalization on novel tasks.
@@
In spatial reasoning tasks, EMMA-X(sun2024emma) exhibits foresighted action planning and fact-based chain-of-thought structures, enabling more effective prediction of environmental changes and corresponding action strategy formulation. SpatialCoT(liu2025spatialcot) employs a two-stage training paradigm of “bidirectional spatial coordinate alignment + CoT spatial alignment” to achieve deep synergy between spatial information and chain-of-thought logic. EmbodiedVSR(zhang2025embodiedvsr) explicitly models structured knowledge via dynamic scene graphs, granting zero-shot spatial reasoning abilities and enhancing adaptability and robustness in unseen environments.
@@
For navigation tasks, MCoCoNav(shen2025enhancing) introduces multimodal chain-of-thought for multi-robot collaborative semantic navigation, designing strategies that integrate global semantic maps with direct inter-robot communication to significantly improve navigation efficiency while reducing communication costs. NavCoT(lin2025navcot) achieves self-guided navigation decision-making at low cost through parameter-efficient domain-specific training, simultaneously enhancing cross-domain generalization capabilities.
@@
Notably, COUT(zhang2024smartagent) pioneers preliminary exploration into learning personalized embodied agents by incorporating individual factors such as user personality and preferences into the autonomous agent’s learning process, thereby broadening embodied intelligence’s application prospects in personalized human–machine interaction.

### v2

@@ 3.4.3 Process Supervision
In the post-training stage, the focus of supervision shifts from the correctness of the outcome to the rationality of the process. Datasets with Chain-of-Thought annotations, such as Visual CoT(Shao et al., 2024a) and LLaVA-CoT-100K(Xu et al., 2024a), provide the model with detailed reasoning paths, enabling it to perform efficient behavioral cloning through supervised fine-tuning or to undergo goal-oriented optimization based on reward signals via reinforcement learning methods.
During the inference stage, the supervision mechanism manifests in a more dynamic and immediate manner. Chain-of-Thought prompting(Wei et al., 2022; Kojima et al., 2022) acts as a non-parametric, on-the-fly form of soft supervision, efficiently guiding the model into a structured reasoning mode. More advanced guidance mechanisms, such as the Process Reward Model proposed by VisVM(Wang et al., 2024g), or the ”critic” role in the multi-agent framework employed by Elhenawy et al.(Elhenawy et al., 2024), constitute external dynamic supervisory signals at inference time. Meanwhile, self-optimization mechanisms, as represented by R3V(Cheng et al., 2024) and EVLM(Khalid et al., 2024), embody a higher-order form of internal dynamic supervision. In this mechanism, the model acts as its own critic, forming an effective closed feedback loop by reflecting on, reviewing, and generating corrective instructions for the intermediate steps it has already produced. Whether originating externally or internally, these dynamic supervisory signals collectively provide fine-grained, real-time calibration of the model’s reasoning behavior.
@@ 4.1 Mathematical and Logical Reasoning
Benchmarks such as MathVista(Lu et al., 2023), MATH-Vision(Wang et al., 2024b), We-Math(Qiao et al., 2024), and CMM-Math(Liu et al., 2024b) integrate visual information into mathematical reasoning tasks. MathVerse(Zhang et al., 2024e) aims to evaluate models’ reasoning abilities on math problems involving charts and adopts a chain-of-thought evaluation strategy, leveraging GPT-4(V) for fine-grained assessment of generated answers. ErrorRadar(Yan et al., 2024) introduces a new multimodal error detection task in a formalized manner, which includes two subtasks: error step identification and error categorization, thus providing a comprehensive framework for assessing complex mathematical reasoning abilities in multimodal large language models.
Additionally, LogicVista(Xiao et al., 2024b) targets the evaluation of comprehensive logical reasoning in visual environments, while MME-Reasoning(Yuan et al., 2025) offers a thorough assessment of logical reasoning capabilities, encompassing inductive, deductive, and abductive reasoning. EnigmaEval(Wang et al., 2025b) explores models’ abilities to integrate implicit knowledge and perform multi-step deductive reasoning by presenting puzzles of varying complexity. MuCR(Li et al., 2024b), a novel multimodal causal reasoning benchmark, leverages synthetic twin image-text pairs to challenge models in identifying causal relationships across different modalities.
@@ 4.2 Spatiotemporal and Directional Reasoning
PulseCheck457(Wang et al., 2025e) offers a comprehensive framework for assessing 6D spatial reasoning across varying levels of complexity. It includes seven question types and spans five difficulty levels, ranging from basic single-object recognition to advanced 6D spatial reasoning tasks. GSR-BENCH(Rajabi & Kosecka, 2024) extends the What’sUp dataset and proposes a novel integrated evaluation method for understanding spatial relations. CDR(Yin et al., 2025) targets the assessment of directional reasoning in multimodal large language models. It includes images annotated with spatial directions (e.g., up, down, left, right) and compass directions (e.g., north, south, east, west), thereby addressing a previously underexplored area in directional reasoning.For temporal reasoning, TOMATO(Shangguan et al., 2024) introduces a new benchmark to assess models’ ability to reason about temporal dynamics in video understanding. In the context of autonomous driving, DriveLMM-o1(Ishaq et al., 2025) provides a benchmark with over 4,000 visual question-answering instances, covering perception, prediction, and planning tasks. Each instance includes a detailed step-by-step reasoning process.
@@ 4.3 Vision-Language Transformation Reasoning
HumanEval-V(Zhang et al., 2024c) employs a novel code generation task to comprehensively evaluate a model’s chart comprehension capabilities. It comprises six task types, each featuring carefully curated charts along with function signatures and test cases. Plot2Code(Wu et al., 2024) provides a comprehensive visual-to-code benchmark that includes high-quality matplotlib charts collected from the matplotlib gallery, accompanied by the corresponding source code and descriptive instructions summarized by GPT-4. Similarly, Plot2XML(Cui et al., 2025b) offers a benchmark for reconstructing scientific plots into editable XML code. P2GB(Chen et al., 2024b) quantitatively evaluates the visual reasoning abilities of multimodal large language models in rich-text or high-resolution image contexts. Current benchmarks for multimodal chain-of-thought (CoT) reasoning predominantly follow the conventional paradigm of multimodal input paired with textual output. To address this limitation, Cheng et al.(Cheng et al., 2025b) proposed CoMT, which requires both multimodal input and multimodal reasoning output, aiming to emulate human behavior in integrating visual operations during the reasoning process.
@@ 4.4 Sequential and Multi-Image Reasoning
MIMU(Zhang et al., 2025b) is developed to assess the performance of multimodal large language models in multi-image input scenarios. It focuses on the model’s ability to perceive fine-grained visual details, with an emphasis on two specific tasks: image-to-image matching and multi-image-to-text matching.
@@ 4.5 Multimodal Integrated Reasoning
MME-CoT(Jiang et al., 2025a), the first comprehensive benchmark for evaluating the CoT reasoning capabilities of multimodal large language models, spans six domains: mathematics, science, optical character recognition (OCR), logic, spatiotemporal reasoning, and general scenarios. M3CoT(Chen et al., 2024c) takes the first step toward addressing multi-domain, multi-step, and multimodal scenarios in multimodal CoT reasoning. Similarly, EMMA(Hao et al., 2025b) and VRC-Bench[84] offer multi-step benchmarks that assess CoT reasoning across modalities and domains. In addition, R1-Onevision-Bench(Yang et al., 2025c) spans multiple subjects—including mathematics, physics, chemistry, biology, and logical reasoning—covering examination content from middle school to university and beyond. NPHardEval4V(Fan et al., 2024) serves as a dynamic benchmark designed to isolate the effects of various factors such as image recognition and instruction following from overall model performance, thereby focusing the evaluation on reasoning capability. Likewise, VisualPuzzles(Song et al., 2025) reduces the reliance on domain-specific knowledge while increasing reasoning complexity to better assess a model’s multimodal reasoning performance.
@@ 4.6 Reasoning Process Quality Evaluation
MPBench(Xu et al., 2025) is a comprehensive multi-task benchmark that systematically evaluates the effectiveness of Process Reward Models (PRMs) through three evaluation paradigms: step correctness identification, answer aggregation, and reasoning path search. MiCEval(Zhou et al., 2024b) assesses the correctness of reasoning chains by evaluating both the quality of model-generated descriptions and each reasoning step. The description evaluation focuses on the accuracy of image descriptions, while the reasoning step evaluation emphasizes the quality of each specific inference. VisualProcessBench(Wang et al., 2025d), a benchmark with manually annotated step-by-step correctness labels, is primarily used to assess the ability of PRMs to detect incorrect reasoning steps in multimodal reasoning tasks. In addition, MMIR(Yan et al., 2025b) and the work by Jia et al.(Jia et al., 2025a) focus on identifying inconsistency issues in multimodal reasoning, offering valuable insights for future improvements of multimodal large language models.
4.7 Evaluation Metrics
At present, most existing works(Xiang et al., 2024a; Thawakar et al., 2025; Li et al., 2024b; Shangguan et al., 2024; Wu et al., 2024) assess the reasoning capabilities of LLMs and MLLMs by directly evaluating the accuracy of their final answers, using this as a proxy for chain-of-thought reasoning performance. Alternatively, some studies propose task-specific evaluation metrics. However, these approaches often suffer from limitations such as a lack of comprehensive evaluation and inability to accurately assess complex visual reasoning tasks. To address these issues, MME-CoT(Jiang et al., 2025a) proposed the first dedicated evaluation framework for assessing CoT reasoning capabilities in MLLMs, which covers three key aspects: CoT quality, CoT robustness, and CoT efficiency. In addition, MMMR(Tie et al., 2025) introduced a modular Reasoning Trace Evaluation Pipeline (RTEP) that provides a holistic assessment of reasoning steps based on three dimensions: relevance to the question, relevance to the answer, and internal consistency across reasoning steps.
@@ 4.7 Evaluation Metrics
MathVista(Lu et al., 2023) 2024 Math QA T, I Accuracy MC, Open 6,141
MATH-Vision(Wang et al., 2024b) 2024 Math QA T, I Accuracy MC, Open 3,040
We-Math(Qiao et al., 2024) 2024 Math QA T, I Custom Metric MC 6.5K
LogicVista(Xiao et al., 2024b) 2024 Logical Reasoning T, I Accuracy MC, Open 448
MathVerse(Zhang et al., 2024e) 2024 Math QA T, I Accuracy, CoT-Evaluation MC, Open 15,672
ErrorRadar(Yan et al., 2024) 2024 Error Detection T, I Accuracy Open 2,500
CMM-Math(Liu et al., 2024b) 2024 Math QA T, I Accuracy MC, Open 28,069
MuCR(Li et al., 2025b) 2025 Causal Reasoning T, I Custom Metric Open ——
EnigmaEval(Wang et al., 2025b) 2025 Puzzle Solving T, I Accuracy Open 1,184
MME-Reasoning(Yuan et al., 2025) 2025 Logical Reasoning T, I Custom Metric MC, Open, Rule-based 1,188
@@
CDR(Yin et al., 2025) 2024 Directional Reasoning T, I Accuracy MC 100K
GSR-BENCH(Rajabi & Kosecka, 2024) 2024 Spatial Reasoning T, I Accuracy MC 4,958
TOMATO(Shangguan et al., 2024) 2024 Temporal Reasoning T, V Custom Metric MC 1,484
PulseCheck457(Wang et al., 2025e) 2025 Spatial Reasoning T, I RPDR Open ——
DriveLMM-o1(Ishaq et al., 2025) 2025 Autonomous Driving T, I, PC Custom Metric MC, Open ——
@@
Plot2Code(Wu et al., 2024) 2024 Visual Coding T, I Custom Metric Open 132
P2G(Chen et al., 2024b) 2024 Visual Reasoning T, I Accuracy MC 2,130
HumanEval-V(Zhang et al., 2024c) 2025 Visual Coding T, I pass@k Open 100K
CoMT(Cheng et al., 2025b) 2025 Multi-task T, I Accuracy, Macro-F1 MC 3,853
Plot2XML(Cui et al., 2025b) 2025 Visual Coding T, I Custom Standard Open 247
@@
MIMU(Zhang et al., 2025b) 2025 Multi-task T, I Accuracy Open ——
@@
NPHardEval4V(Fan et al., 2024) 2024 VQA T, I Custom Metric Open ——
M3CoT(Chen et al., 2024c) 2024 VQA T, I Accuracy MC 11,459
EMMA(Hao et al., 2025b) 2025 VQA T, I pass@k MC, Open 2,788
VRC-Bench(Thawakar et al., 2025) 2025 Multi-task T, I Custom Multi-metric MC, Open ——
MME-CoT(Jiang et al., 2025a) 2025 VQA T, I Quality, Robustness, Efficiency MC, Open 1,130
R1-Onevision-Bench(Yang et al., 2025c) 2025 VQA T, I Accuracy MC, Open 942
VisualPuzzles(Song et al., 2025) 2025 VQA T, I Accuracy MC 1,168
@@
MiCEval(Zhou et al., 2024b) 2025 VQA T, I Custom Metric Open 2,130
MMIR(Yan et al., 2025b) 2025 Consistency Reasoning T, I Accuracy MC, Open 534
VisualProcessBench(Wang et al., 2025d) 2025 Science QA T, I Macro-F1 MC, Open 2,866
@@ 5.1 Embodied AI
In the domain of task decomposition and execution, the works by Hao et al.(Hao et al., 2025a) and Embodied Reasoner(Zhang et al., 2025e) demonstrate the feasibility of automatically dividing operation instructions or task goals into concrete, detailed execution steps. Specifically, Hao et al. enhance humanoid robots’ abilities to perform localized operations in unstructured complex scenarios by incorporating foundation models, while Embodied Reasoner targets interactive search tasks, showcasing the agent’s capacity for autonomous goal recognition, reasoning, and decision-making, thereby strengthening adaptability and proactivity in dynamic tasks.
@@
Regarding task planning and hierarchical coordination, EmbodiedGPT(Mu et al., 2023) proposes a unified architecture that integrates high-level strategic planning with low-level control execution, enabling embodied agents to accomplish complex tasks through stepwise planning. The MCoT-Memory framework(Liang et al., 2024) introduces a memory retrieval–based, scene graph–driven chain-of-thought mechanism that retains high-confidence experiential knowledge to effectively support long-term and complex task planning in dynamic environments. ECoT(Zawalski et al., 2024) further expands embodied intelligence’s multimodal reasoning capabilities by incorporating chain-of-thought training mechanisms into Vision-Language-Action (VLA) models, allowing for multi-step task reasoning and subtask decomposition with strong generalization on novel tasks.
@@
In spatial reasoning tasks, EMMA-X(Sun et al., 2024c) exhibits foresighted action planning and fact-based chain-of-thought structures, enabling more effective prediction of environmental changes and corresponding action strategy formulation. SpatialCoT(Liu et al., 2025b) employs a two-stage training paradigm of “bidirectional spatial coordinate alignment + CoT spatial alignment” to achieve deep synergy between spatial information and chain-of-thought logic. EmbodiedVSR(Zhang et al., 2025f) explicitly models structured knowledge via dynamic scene graphs, granting zero-shot spatial reasoning abilities and enhancing adaptability and robustness in unseen environments.
@@
For navigation tasks, MCoCoNav(Shen et al., 2025b) introduces multimodal chain-of-thought for multi-robot collaborative semantic navigation, designing strategies that integrate global semantic maps with direct inter-robot communication to significantly improve navigation efficiency while reducing communication costs. NavCoT(Lin et al., 2025a) achieves self-guided navigation decision-making at low cost through parameter-efficient domain-specific training, simultaneously enhancing cross-domain generalization capabilities.
@@
Notably, COUT(Zhang et al., 2024d) pioneers preliminary exploration into learning personalized embodied agents by incorporating individual factors such as user personality and preferences into the autonomous agent’s learning process, thereby broadening embodied intelligence’s application prospects in personalized human–machine interaction.